OpenStudy (anonymous):

The volume of a cone is 62.8 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? Use π = 3.14.

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Rushwr

OpenStudy (rushwr):

\[V= \Pi r ^{3}\frac{ h }{ 3 }\]

OpenStudy (rushwr):

This is the volume of the cone . We just have to substitute the numbers here !

OpenStudy (anonymous):

ok!

OpenStudy (welshfella):

Volume of a cone = (1/3) pi r^2 h where r = radius and h = height so plug in the given values 62.8 = (1/3) * 3.14 * r^2 * 15 solve for r

OpenStudy (rushwr):

Since they have asked us the radius in inches no need to convert the height and the radius in to meters.

OpenStudy (welshfella):

@Rushwr - is that a typo ? r^2 not r^3

OpenStudy (anonymous):

ok i get it just trying to figure it out

OpenStudy (rushwr):

oh pellet sorry

OpenStudy (rushwr):

human error kid ! it is r^2

OpenStudy (rushwr):

Sorry abwt tht! I actually thought I typd 2

OpenStudy (welshfella):

don't worry we all do it sometimes...

OpenStudy (rushwr):

So at the end the answer would be 2 inches !

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

alright i got the answer thanks a lot guysss

OpenStudy (welshfella):

yw